OpenStudy (anonymous):

solve by he elimination method 5r-7s=-7 7r+5s=79

OpenStudy (anonymous):

r= (-7 + 7s) / 5 Plug that into 2nd equation. [(-49 +49s) /5] +5s =79 -49 + 74s = 395 Therefore s = 6

OpenStudy (anonymous):

plug s=6 into 1st now. 5r = 35 Therefore r = 7

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

rate me xD